In Triangle ABC, angle C = 90°, then find the value of sec (A+B) 

Grade:10

3 Answers

Arun
4 years ago
Dear student
 
sec (A + B) = sec ( 90) = not defined.
 
Hope it helps
 
Regards
Arun
 
In case of any difficulty, please feel free to ask again 
Vikas TU
4 years ago
Dear student 
m(angle A) + m(angle B ) + m (angleC ) = 180 °…..sum of the three angles of a triangle is 180°
m(angle A) + m(angle B ) + 90° = 180°
Where angle C is 90 ° (C is right angle given)
m(angle A + angle B ) = 180°- 90° = 90°
Cos( A+ B) = 90°
Cos 90 ° = 0
Sec (A+B) = 1/0 = not defined.
Hope this helps 
Good Luck 
sushmitha raj
4 years ago
angleA+angleB+angleC=180o
given,
C=90o
A+B+90=180
A+B=90
Cos(A+B)=cos90
cos(A+B)=0
Sec(A+B)=1/0
Sec(A+B)=not defined